Royster Company presents the following income statement:

Sales $12,000
Cost of goods sold $6,000
Selling and administrative expense $1,200
Interest expense $600
Pretax income $4,200
Income tax expense $1,470
Net income $2,730


Which of the following line items would appear on a common-size income statement for this period?

Net income 65%

Income tax expense 54%

Pretax income 35%
Answer #1

Each item such as expense,pretax income,net income etc. is represented as a % of net sales in common-size income statement.

Hence net income%=(2730/12000)=22.75%

Income tax expense%=(1470/12000)=12.25%

Pretax income%=(4200/12000)=35%

Hence the correct option is C.